# ChronoRoot 2.0: An Open AI-Powered Platform for 2D Temporal Plant Phenotyping

### v1

@@ Abstract
The analysis of plant developmental plasticity, including root system architecture, is fundamental to understanding plant adaptability and development, particularly in the context of climate change and agricultural sustainability. While significant advances have been made in plant phenotyping technologies, comprehensive temporal analysis of root development remains challenging, with most existing solutions providing either limited throughput or restricted structural analysis capabilities. Here, we present ChronoRoot 2.0, an integrated open-source platform that combines affordable hardware with advanced artificial intelligence to enable sophisticated temporal plant phenotyping. The system introduces several major advances, offering an integral perspective of seedling development: (i) simultaneous multi-organ tracking of six distinct plant structures, (ii) quality control through real-time validation, (iii) comprehensive architectural measurements including novel gravitropic response parameters, and (iv) dual specialized user interfaces for both architectural analysis and high-throughput screening. We demonstrate the system’s capabilities through three use cases for Arabidopsis thaliana: characterization of circadian growth patterns under different light conditions, detailed analysis of gravitropic responses in transgenic plants, and high-throughput screening of etiolation responses across multiple genotypes. ChronoRoot 2.0 maintains its predecessor’s advantages of low cost and modularity while significantly expanding its capabilities, making sophisticated temporal phenotyping more accessible to the broader plant science community. The system’s open-source nature, combined with extensive documentation and containerized deployment options, ensures reproducibility and enables community-driven development of new analytical capabilities.
@@ 1 Introduction
Building upon this foundation, we present ChronoRoot 2.0, which significantly expands the capabilities and accessibility of the platform through several key innovations, as illustrated in the comprehensive pipeline shown in Figure 1. The complete workflow begins with the Hardware Module (Fig. 1-01) that combines Raspberry Pi-controlled cameras and infrared LED backlighting. This setup enables consistent Image Acquisition (Fig. 1-02) irrespective of day/night cycles, with a temporal resolution of 15 minutes for extended monitoring of plant development.
@@
At the core of these improvements is an upgraded nnUNet Segmentation module (Fig. 1-03) based on the nnUNet architecture [7], which performs simultaneous multi-class segmentation of six distinct plant structures: main root (class 1), lateral roots (class 2), seed (class 3), hypocotyl (class 4), leaves (class 5) and petiole (class 6). This advancement enables comprehensive tracking of plant development from seed to mature seedling, capturing the intricate relationships between different organs during growth. The choice of nnUNet as the core architecture was motivated by its proven success in medical image segmentation and its ability to self-configure hyperparameters, making it particularly suitable for biologists without extensive machine learning expertise. While the segmentation model is trained to work with Arabidopsis thaliana, it could easily be retrained for other species. This architectural choice, combined with newly developed graphical user interfaces, significantly lowers the barrier to entry for researchers seeking to implement plant phenotyping in their work.
@@
After segmentation, researchers can select between two distinct but complementary interfaces based on their experimental needs. The Standard Root Phenotyping Interface (Fig. 1-04) maintains continuity with the original ChronoRoot design, focusing on detailed RSA analysis of individual plants. This interface provides researchers with tools for precise measurement and analysis of root development patterns through a graph-based representation approach. It maintains the core strengths of the original ChronoRoot system while adding comprehensive visualization capabilities and an intuitive graphical user interface for analyzing basic architecture, growth dynamics, spatial distribution, and newly determined angular measurements.
@@
The new Screening Interface (Fig. 1-05) extends the system’s capabilities by enabling automated analysis of multiple plants simultaneously. It incorporates the Simple Online Realtime Tracking (SORT) algorithm[8] for robust plant identification across experimental groups, along with manual calibration tools for standardized measurements. This interface specializes in early development analysis through three dedicated modules: germination analysis, hypocotyl analysis, and plant analysis, enabling researchers to efficiently process and compare multiple experimental conditions.
@@ 2.2 Hardware Implementation and Image Acquisition
The four-plate module is compact (62 x 36 x 20 cm) and can be placed in any standard plant growth chamber. The different parts of the imaging setup (back-light, plate support and camera) can be positioned along a horizontal double-rail to control the field of view of the camera and accurate lighting. In addition, the camera can be moved vertically. ChronoRoot allows image acquisition at a high temporal resolution (a set of pictures every minute). The use of an IR back-light (850 nm) and optional long pass IR filters ($>$ 830 nm) allow acquiring images of the same quality independently from the light conditions required for the experiment, during day and night.
@@ 2.3 Deep Learning Segmentation Framework
The foundation of ChronoRoot 2.0’s machine learning capabilities rests on a comprehensive dataset of Arabidopsis developmental sequences. This dataset represents a significant expansion from the original ChronoRoot, encompassing over 693 manually annotated images that capture multiple plants across various developmental stages. Expert biologists performed detailed annotation of seven distinct structural classes using ITK-SNAP [9], chosen for its precise annotation tools and user-friendly interface.
@@
Image acquisition utilized the ChronoRoot hardware system’s standardized configuration, producing infrared images at 3280 x 2464 resolution. The single-channel nature of infrared imaging enables efficient data processing while maintaining all relevant structural information. To ensure robust model generalization, the dataset incorporates specimens from various experimental conditions and genetic backgrounds.
The segmentation framework implements the nnUNet architecture [7], selected for its self-configuring capabilities and proven performance in biomedical image analysis. The network dynamically adjusts its depth, width, kernel sizes, stride patterns, normalization schemes, and learning rate schedules based on the training dataset’s properties, eliminating the need for extensive hyperparameter tuning. This adaptability is particularly valuable in plant phenotyping contexts, where morphological diversity and varying experimental conditions create unique imaging challenges.
@@ 2.5.1 Standard Root System Architecture Analysis
The analysis begins with user definition of Regions of Interest (ROIs), allowing precise selection of individual plants from multi-plant images. Within these ROIs, the system processes segmentation masks through morphological operations and thinning algorithms to obtain skeletal representations of the root system. These skeletons are then analyzed to identify key nodes, which serve as the basis for constructing a graph representation using a depth-first search (DFS) algorithm [12].
@@
Figure 2: Angular measurements in root system architecture. Illustration of base-tip angle ($\theta_{bt}$) and emergence angle ($\theta_{e}$) calculations showing how these complementary metrics quantify different aspects of lateral root orientation.
@@
Basic Architecture Main Root (MR) Length mm Path length along graph skeleton 1
Lateral Root (LR) Length mm Sum of all LR path lengths 1
Total Root (TR) Length mm MR Length + LR Length 1
@@
Fourier Components Hz FFT of detrended time series 1
@@
Convex Hull Width mm Maximum horizontal extent 2
Convex Hull Height mm Maximum vertical extent 2
@@
Hypocotyl Length mm Path length of hypocotyl skeleton 3
@@
Simple Root Length mm Path length of primary root skeleton 3
@@ 2.5.2 Multiple Plant Screening Analysis
The hypocotyl analysis module incorporates validation steps for reliable measurement in multi-plant scenarios. The system automatically detects physiologically impossible growth rates and artifacts that can occur when different plants touch and their segmentations combine. Biological constraints are enforced through validation of non-decreasing length measurements and growth directions against expected gravitropic responses.
@@ 2.6 Software Implementation and User Interface
The Standard Root Phenotyping Interface maintains the core functionality of the original ChronoRoot system while adding modern visualization capabilities. This interface implements a comprehensive analysis pipeline through several interconnected modules (Supplementary Figure S2). The main interface provides tools for experimental configuration, ROI-based plant selection, and real-time visualization of segmentation results. Through an intuitive workflow, users can configure analysis parameters, process individual plants, and generate detailed architectural measurements. The analysis capabilities include convex hull analysis, lateral root angle measurements, growth speed evaluation with Fourier analysis, and detailed statistical testing using Mann-Whitney tests at configurable time intervals. Users can specify particular days for detailed reporting and adjust various measurement parameters such as emergence distance for lateral roots. The interface includes quality control through visual feedback systems and supports automated report generation with customizable visualization options.
@@ 3 Results
The performance and capabilities of ChronoRoot 2.0 were evaluated through three key aspects. First, we assessed the core segmentation capabilities, comparing our nnUNet implementation against the original ChronoRoot system in both accuracy and computational efficiency. Second, we validated the system’s multi-class detection capabilities, evaluating its performance in simultaneously identifying and tracking six distinct plant structures. Finally, we demonstrated the software’s practical utility through three comprehensive use cases, showcasing its application in both detailed architectural analysis and high-throughput screening scenarios.
@@ 3.1 Segmentation Performance with nnUNet
Table 2: Segmentation performance comparison between original ChronoRoot and ChronoRoot 2.0. The nnUNet implementation outperforms previous models in both accuracy (Dice coefficient) and boundary precision (Hausdorff distance), with individual folds surpassing the original ensemble approach. The final nnUNet ensemble achieves 6% higher Dice score and 19% lower Hausdorff distance than the original ensemble.
Model Dice Hausdorff (mm) Processing Time
@@
DSResUNet (Fast) 0.769 $\pm$ 0.048 8.14 $\pm$ 7.34 $\sim 0.5s$
Ensemble (Accurate) 0.772 $\pm$ 0.048 6.68 $\pm$ 5.08 $\sim 4.5s$
@@
Fold 0 0.817 $\pm$ 0.030 5.58 $\pm$ 5.52 $\sim 3s$
Fold 1 0.811 $\pm$ 0.035 5.65 $\pm$ 5.40 $\sim 3s$
Fold 2 0.813 $\pm$ 0.029 5.60 $\pm$ 5.48 $\sim 3s$
Fold 3 0.813 $\pm$ 0.033 4.74 $\pm$ 4.81 $\sim 3s$
Fold 4 0.814 $\pm$ 0.030 6.06 $\pm$ 7.36 $\sim 3s$
nnUNet Ensemble 0.817 $\pm$ 0.029 5.38 $\pm$ 5.53 $\sim 15s$
We first evaluated the segmentation performance of ChronoRoot 2.0’s nnUNet implementation against the original ChronoRoot models using their established dataset and metrics, to validate our architectural improvements. This comparison not only validates the new segmentation approach but also demonstrates backward compatibility with the original system’s binary segmentation task, ensuring continuity for existing users while providing enhanced capabilities. The nnUNet implementation showed substantial accuracy gains while maintaining practical processing speeds for high-throughput applications (Table 2).
@@
Regarding ChronoRoot 2.0, since the nnUNet consists of an ensemble of 5 different networks trained with different sub-folds generated from the training data, we evaluated both, the performance of the individual networks (Fold 0 to 4) and the nnUNet Ensemble. We found that nnUNet exhibits an optimal balance between accuracy and speed. Individual networks consistently outperformed the original ensemble method, achieving Dice scores above 0.81 with processing times of  3 seconds per image. The full ensemble further improved accuracy (Dice: 0.817 ± 0.029) while reducing structural misalignment (Hausdorff distance: 5.38 ± 5.53 mm). These improvements enable more reliable architectural measurements while maintaining throughput suitable for time-series experiments.

### v2

@@ Abstract
Background: Plant developmental plasticity, particularly in root system architecture, is fundamental to understanding adaptability and agricultural sustainability. Existing automated phenotyping solutions face limitations including binary segmentation approaches, restricted structural analysis capabilities, and text-based interfaces that limit accessibility, with most focusing solely on root structures while overlooking valuable information from simultaneous analysis of multiple plant organs.
Findings: ChronoRoot 2.0 builds upon established low-cost hardware while significantly enhancing software capabilities and usability. The system employs nnUNet architecture for multi-class segmentation, demonstrating significant accuracy improvements while simultaneously tracking six distinct plant structures encompassing root, shoot, and seed components: main root, lateral roots, seed, hypocotyl, leaves, and petiole. This architecture enables easy retraining and incorporation of additional training data without requiring machine learning expertise. The platform introduces dual specialized graphical interfaces: a Standard Interface for detailed architectural analysis with novel gravitropic response parameters, and a Screening Interface enabling high-throughput analysis of multiple plants through automated tracking. Functional Principal Component Analysis integration enables discovery of novel phenotypic parameters through temporal pattern comparison. We demonstrate multi-species analysis, with Arabidopsis thaliana and Solanum lycopersicum, both morphologically distinct plant species. Three use cases in Arabidopsis thaliana and validation with tomato seedlings demonstrate enhanced capabilities: circadian growth pattern characterization, gravitropic response analysis in transgenic plants, and high-throughput etiolation screening across multiple genotypes.
Conclusions: ChronoRoot 2.0 maintains the low-cost, modular hardware advantages of its predecessor while dramatically improving accessibility through intuitive graphical interfaces and expanded analytical capabilities. The open-source platform makes sophisticated temporal plant phenotyping more accessible to researchers without computational expertise.
Software availability: https://chronoroot.github.io
keywords:
Plant phenotyping; Root system architecture; Deep learning segmentation; Temporal analysis; High-throughput screening; Open-source software; Arabidopsis thaliana; Tomato
@@ 1 Introduction
Building upon this foundation, we present ChronoRoot 2.0, which significantly expands the capabilities and accessibility of the platform through several key innovations, as illustrated in the comprehensive pipeline shown in Figure 1. The complete workflow begins with the Hardware Module (Fig. 1-A) that combines Raspberry Pi-controlled cameras and infrared LED backlighting. This setup enables consistent Image Acquisition (Fig. 1-B) irrespective of day/night cycles, with a temporal resolution of 15 minutes for extended monitoring of plant development.
@@
At the core of these improvements is an upgraded nnUNet Segmentation module (Fig. 1-C) based on the nnUNet architecture [7], which performs simultaneous multi-class segmentation of six distinct plant structures: main root (class 1), lateral roots (class 2), seed (class 3), hypocotyl (class 4), leaves (class 5) and petiole (class 6). This advancement enables comprehensive tracking of plant development from seed to mature seedling, capturing the intricate relationships between different organs during growth. The choice of nnUNet as the core architecture was motivated by its proven success in medical image segmentation and its ability to self-configure hyperparameters, making it particularly suitable for biologists without extensive machine learning expertise. While the segmentation model is trained to work with both Arabidopsis thaliana and Solanum lycopersicum (tomato), it can be easily adapted for other species, leveraging the self-configuring capabilities of the nnUNet architecture. This architectural choice, combined with newly developed graphical user interfaces, significantly lowers the barrier to entry for researchers seeking to implement plant phenotyping in their work.
@@
After segmentation, researchers can select between two distinct but complementary interfaces based on their experimental needs. The Standard Root Phenotyping Interface (Fig. 1-D) maintains continuity with the original ChronoRoot design, focusing on detailed RSA analysis of individual plants. This interface provides researchers with tools for precise measurement and analysis of root development patterns through a graph-based representation approach. It maintains the core strengths of the original ChronoRoot system while adding comprehensive visualization capabilities and an intuitive graphical user interface for analyzing basic architecture, growth dynamics, spatial distribution, and newly determined angular measurements.
@@
The new Screening Interface (Fig. 1-E) extends the system’s capabilities by enabling automated analysis of multiple plants simultaneously. It incorporates the Simple Online Realtime Tracking (SORT) algorithm[8] for robust plant identification across experimental groups, along with manual calibration tools for standardized measurements. This interface specializes in early development analysis through three dedicated modules: germination analysis, hypocotyl analysis, and plant analysis, enabling researchers to efficiently process and compare multiple experimental conditions.
@@ 2.2 Hardware Implementation and Image Acquisition
The four-plate module is compact (62 x 36 x 20 cm) and can be placed in any standard plant growth chamber. The different parts of the imaging setup (back-light, plate support and camera) can be positioned along a horizontal double-rail to control the field of view of the camera and accurate lighting. In addition, the camera can be moved vertically. ChronoRoot allows image acquisition at a high temporal resolution (a set of pictures every minute). The use of an IR back-light (850 nm) and optional long pass IR filters (>> 830 nm) allow acquiring images of the same quality independently from the light conditions required for the experiment, during day and night.
@@ 2.3 Deep Learning Segmentation Framework
ChronoRoot 2.0’s machine learning capabilities rests on a comprehensive dataset of plant developmental sequences that represents a significant expansion from the original ChronoRoot. The dataset comprises 911 manually annotated images of Arabidopsis thaliana and 480 images of tomato, capturing multiple plants across various developmental stages. Expert biologists performed detailed annotation using ITK-SNAP [9], chosen for its precise annotation tools and user-friendly interface.
@@
In the tomato dataset, leaves and petiole were annotated as a single combined aerial part class, reflecting species-specific morphological differences.
Image acquisition utilized the ChronoRoot hardware system’s Raspberry Pi Noir V2 camera, producing infrared images at 3280 x 2464 resolution (approximately $\approx$ 0.04 mm per pixel). Images are processed at single channel full native resolution, preserving all structural detail throughout both training and inference. To ensure robust model generalization, the dataset incorporates specimens from various experimental conditions and genetic backgrounds.
The segmentation framework implements the nnUNet architecture [7], selected for its self-configuring capabilities and proven performance in biomedical image analysis. The network dynamically adjusts its depth, width, kernel sizes, stride patterns, normalization schemes, and learning rate schedules based on the training dataset’s properties, eliminating the need for extensive hyperparameter tuning. This adaptability is particularly valuable in plant phenotyping contexts, where morphological diversity and varying experimental conditions create unique imaging challenges. Our implementation uses the official nnUNet v2 framework, facilitating straightforward adaptation to new plant species through standard retraining procedures. The framework provides two architectural variants: the standard convolutional nnUNet and the recently introduced nnUNet with residual encoder connections [10], which improves feature learning through skip connections in the encoder pathway. Additionally, the framework supports optional test-time augmentation, where predictions from multiple augmented versions of each image are averaged to improve segmentation robustness, particularly for boundary detection.
@@ 2.5.1 Standard Root System Architecture Analysis
The analysis begins with user definition of ROIs, allowing precise selection of individual plants from multi-plant images. Within these ROIs, the system processes segmentation masks through morphological operations and thinning algorithms to obtain skeletal representations of the root system. These skeletons are then analyzed to identify key nodes, which serve as the basis for constructing a graph representation using a depth-first search algorithm [13].
@@
Figure 2: Angular measurements in root system architecture. Illustration of base-tip angle ($\theta_{bt}$) and emergence angle ($\theta_{e}$) calculations on a Arabidopsis thaliana plant showing how these complementary metrics quantify different aspects of lateral root orientation.
@@
To prepare the data for final analysis and visualization, all measurements pass through an automated post-processing script that ensures biological and structural consistency across the time series. To account for skeletonization noise and avoid the inclusion of false-positive "spurs" on the root axes, we implemented a structural pruning threshold where a minimum length of 5 pixels ($\approx$ 0.2 mm) is required for a skeleton branch to be preserved prior to graph construction. Additionally, a measurement post-processing script filters out "false starts" and transient misdetections through a temporal verification window: any detected structure is only validated if it persists beyond a 6-hour threshold. Finally, the pipeline enforces a monotonic growth constraint for length measurements, preventing impossible decreases in plant size over time. While these cleaned results are used for the study figures, the raw segmentations, graphs, and original RSML data remain available for researchers who wish to perform their own specific analyses.
@@
Basic Architecture Main Root (MR) Length $m$ Path length along graph skeleton 1
Lateral Root (LR) Length $m$ Sum of all LR path lengths 1
Total Root (TR) Length $m$ MR Length + LR Length 1
@@
Fourier Components Hz Fast Fourier Transform of detrended time series 1
@@
Convex Hull Width $m$ Maximum horizontal extent 2
Convex Hull Height $m$ Maximum vertical extent 2
@@
Hypocotyl Length $m$ Path length of hypocotyl skeleton 3
@@
Simple Root Length $m$ Path length of primary root skeleton 3
@@ 2.5.2 Multiple Plant Screening Analysis
The hypocotyl analysis module incorporates validation steps for reliable measurement in multi-plant scenarios. The system automatically detects physiologically impossible growth rates and artifacts that can occur when different plants touch and their segmentations combine, incorporating biological constraints by forcing non-decreasing length measurements. To ensure full transparency, the platform preserves all raw, unfiltered measurements alongside the processed results. These files are exported in standard formats, allowing researchers to apply custom validation logic or modify the underlying code for specialized needs.
@@ 2.6 Software Implementation and User Interface
The Standard Root Phenotyping Interface maintains the core functionality of the original ChronoRoot system while adding modern visualization capabilities. This interface implements a comprehensive analysis pipeline through several interconnected modules (Supplementary Figure S2). The main interface provides tools for experimental configuration, ROI-based plant selection, and real-time visualization of segmentation results. Through an intuitive workflow, users can configure analysis parameters, process individual plants, and generate detailed architectural measurements. The analysis capabilities include convex hull analysis, lateral root angle measurements, growth speed evaluation with Fourier analysis, and detailed statistical testing using Mann-Whitney tests at configurable time intervals. Users can specify particular days for detailed reporting and adjust various measurement parameters such as emergence distance for lateral roots. The interface incorporates quality control through visual feedback systems that allow users to inspect the segmentation results prior to plant selection and manually define the root starting position. Once processed, the software generates growth videos overlaid with the resulting graphs and showcases the measurements, enabling researchers to visually validate the tracking performance for each plant. Problematic individuals can then be discarded or re-analyzed before the system proceeds to the automated generation of comprehensive reports and statistical summaries.
@@ 3 Results
The performance and capabilities of ChronoRoot 2.0 were evaluated through four key aspects. First, we assessed the core segmentation capabilities, comparing our nnUNet implementation against the original ChronoRoot system in both accuracy and computational efficiency. Second, we validated the system’s multi-class detection capabilities, evaluating its performance in simultaneously identifying and tracking six distinct plant structures across both Arabidopsis thaliana and tomato. Third, we demonstrated the system’s multi-species capability through comprehensive evaluation on both species, showcasing robust performance across morphologically distinct plants. Finally, we demonstrated the software’s practical utility through four comprehensive use cases, showcasing its application in both detailed architectural analysis and high-throughput screening scenarios.
@@ 3.1 Segmentation Performance with nnUNet
Table 2: Segmentation performance comparison between original ChronoRoot and ChronoRoot 2.0. The nnUNet implementation outperforms previous models in both accuracy and processing speed in a separated test set (n=55). All nnUNet configurations achieve higher Dice scores than the original models. While test-time augmentation (TTA) shows no significant impact on segmentation overlap (Dice), it substantially improves boundary precision (Hausdorff), reducing error distances by removing spurious segmentations.
Model Dice Hausdorff (mm) Processing Time (s)
@@
DSResUNet (Fast) 0.769 $\pm$ 0.043 7.25 $\pm$ 6.87 $\sim$0.5
Ensemble (Accurate) 0.772 $\pm$ 0.048 7.21 $\pm$ 7.02 $\sim$4.5
@@
Standard 0.809 $\pm$ 0.041 6.41 $\pm$ 5.39 2.80 $\pm$ 0.09
Standard (no TTA) 0.808 $\pm$ 0.042 11.08 $\pm$ 13.58 0.89 $\pm$ 0.07
Residual 0.812 $\pm$ 0.038 9.07 $\pm$ 9.33 5.34 $\pm$ 0.15
Residual (no TTA) 0.815 $\pm$ 0.032 13.10 $\pm$ 14.18 1.57 $\pm$ 0.04
We first evaluated the segmentation performance of ChronoRoot 2.0’s nnUNet implementation against the original ChronoRoot models using their established dataset (consisting of 339 train images and 55 test images) and metrics (Dice coefficient quantifies the overlap between predicted and ground truth segmentations, while the Hausdorff distance measures the maximum boundary error in millimeters), to validate our architectural improvements. This comparison not only validates the new segmentation approach but also demonstrates backward compatibility with the original system’s binary segmentation task, ensuring continuity for existing users while providing enhanced capabilities. The nnUNet implementation showed substantial accuracy gains while maintaining practical processing speeds for high-throughput applications (Table 2).
@@
Regarding ChronoRoot 2.0, we trained two different nnUNet architectural configurations: the standard convolutional architecture and a novel incorporation of a residual encoder architecture [10]. Our implementation allows users to activate or deactivate test-time augmentation (TTA) at inference time, providing a flexible trade-off between processing speed and segmentation quality. All nnUNet configurations substantially outperformed the original ChronoRoot models, achieving Dice coefficients above 0.808 while maintaining practical processing speeds. Disabling TTA reduces inference time by approximately 3-fold (from 2.80 to 0.89 seconds for standard architecture, and from 5.34 to 1.57 seconds for residual), enabling high-throughput processing. Importantly, test-time augmentation showed divergent effects on the two evaluation metrics: TTA had no significant impact on Dice coefficients, comparing architectures with and without TTA revealed nearly identical overlap performance, yet dramatically improved boundary precision as measured by Hausdorff distance, reducing error distances by 40-45%. This improvement stems from TTA’s ability to remove spurious segmentations through prediction averaging, which primarily affects boundary outliers rather than overall segmentation overlap.
@@
Table 3: Multi-class segmentation performance on the Arabidopsis thaliana dataset. All model configurations achieve similar results (n=176). Notably, the fast variants provide a significant reduction in processing time with a minor loss in segmentation performance or structural correctness.
Model Dice HD Cp Cr Time (s)
Standard 0.763±0.196 8.519±12.500 0.934±0.074 0.937±0.106 2.821
Standard (Fast) 0.758±0.198 9.199±13.312 0.929±0.082 0.936±0.110 0.972
Residual 0.764±0.193 8.415±11.983 0.930±0.104 0.937±0.109 5.300
Residual (Fast) 0.763±0.189 8.743±12.324 0.926±0.109 0.935±0.113 1.630